OpenStudy (anonymous):

What is the equation of the line that passes through (0, -2) and has a slope of 0? y = -2 y = 2 x = 2

OpenStudy (anonymous):

Plug it into point-slope form

OpenStudy (anonymous):

\(\sf y - y_1 = m(x - x_1)\)

OpenStudy (anonymous):

(0, -2) x1 y1 m = 0

OpenStudy (anonymous):

Or point-intercept y=mx+b

OpenStudy (anonymous):

since the y intercept is -2 ....

OpenStudy (anonymous):

\(\sf y + 2 = 0(x - 0)\)

OpenStudy (anonymous):

Simplify

OpenStudy (anonymous):

The right side equals 0

OpenStudy (anonymous):

y + 2 = 0

OpenStudy (anonymous):

Subtract 2 to both sides

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

Yes, so we have: y = -2